OpenStudy (anonymous):

A software company is planning for an upgrade of their software. You must charge customers $100. Are your customers willing to pay this much? You contact a random sample of 40 customers and find that 11 would pay $100 for the upgrade. If the upgrade is to be profitable, you will need to sell it to more than 20% of your customers. Do the sample data give good evidence that more than 20% are willing to buy? a. Formulate this problem as a hypothesis test. Give the null and alternative hypotheses. b. Carry out the significance test. Report the test statistic and the P-value.

OpenStudy (anonymous):

c. Should you proceed with plans to produce and market the upgrade?

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

what will you use as your hypothesises?

OpenStudy (amistre64):

you will need to sell it to more than 20% of your customers. Ho is by convention a statement of equality: Ho: p=.20 Ha defines the alternative outcome: Ha: p>.20 seems fair agreed?

OpenStudy (anonymous):

Yes, and b?

OpenStudy (amistre64):

carry out the tests of course ...

OpenStudy (anonymous):

Could you talk me through it?

OpenStudy (amistre64):

\[z=\frac{p-.20}{\sqrt{pq/n}}\]sounds familiar, for some sample proportion.

OpenStudy (amistre64):

what is your sample proportion?

OpenStudy (anonymous):

I'm not sure, I tried to figure it out, but can't seem to get it right.

OpenStudy (amistre64):

how many people were sampled?

OpenStudy (anonymous):

40

OpenStudy (amistre64):

and how many people said they would pay?

OpenStudy (anonymous):

11

OpenStudy (amistre64):

then ... what proportion of the sample said yes?

OpenStudy (anonymous):

.275?

OpenStudy (amistre64):

11/40 = .275 correct, and this means the opposing proportion is? 1- .275 right?

OpenStudy (anonymous):

right

OpenStudy (amistre64):

we are ready to compute the z:\[z=\frac{.275-.20}{\sqrt{\frac{.275(.725)}{40}}}\]

OpenStudy (anonymous):

z=1.062321348?

OpenStudy (amistre64):

hold on, i think i got some things backwards: we use the expected value of .20 underneath

OpenStudy (amistre64):

\[z=\frac{.275-.20}{\sqrt{\frac{.20(.80)}{40}}}\] that is what we want ..... underneath it

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=%5Cfrac%7B.275-.20%7D%7B%5Csqrt%7B%5Cfrac%7B.20%28.80%29%7D%7B40%7D%7D%7D i still get someting less than 1%: .0047

OpenStudy (anonymous):

didn't your equation show that (.275-.20) was over the square root? On the Wolfram page it shows it beside the square root.

OpenStudy (amistre64):

\[\frac{3-2}{5}=\frac{3-1}{1}*\frac{1}{5}=(3-2)*(\frac{1}{5})\] dont fret yourself over multiplication ....

OpenStudy (anonymous):

Thats not what it looked like, could you please humor me and go to the link and look at it because I want to make sure it is correct.

OpenStudy (amistre64):

fine fine fine ... the wolf is spose to read latex but i see that its a little confused :)

OpenStudy (anonymous):

P=0.117839956714519?

OpenStudy (amistre64):

the right tailed area from 1.185 is the P-value

OpenStudy (amistre64):

0.117841... so yes

OpenStudy (amistre64):

as an overview: we are looking to compare our sample value with a normal distribution having a mean of .20 and a standard deviation or sqrt(.20(.80)/40) if that helps visualize this with prior work

OpenStudy (anonymous):

Thanks that helped a bunch can you help me with another question?

OpenStudy (amistre64):

i dont read a significance value in the information, so we tend to be free to use whatever we want. a good rule of thumb is 5%, since 5% of the data is said to occur rarely.

OpenStudy (amistre64):

i can try ;)

OpenStudy (anonymous):

For a single proportion the margin of error of a confidence interval is largest for any given sample size n and confidence level C when p-hat = 0.5. This led us to use p-hat = 0.5 for planning purposes. Use these conservative values in the following calculations, and assume that the sample sizes of n. Calculate the margins of error of the 99% confidence intervals for the following choices of n: 10, 30, 50, 100, 200, and 500. Present the results in a table. Summarize your conclusions.

OpenStudy (anonymous):

I'm not sure how to even approach this one.

OpenStudy (amistre64):

it amounts to reworking the z formula; given that the zscore is equal to the 99% interval mark

OpenStudy (amistre64):

let x be the bound for the z value such that:\[\pm z_{.005}=\frac{E-\mu}{\sigma/\sqrt n}\] solve for E

OpenStudy (anonymous):

How?:}

OpenStudy (amistre64):

algebra of course

OpenStudy (amistre64):

\[\pm z_{.005}=\frac{E-\mu}{\sigma/\sqrt n}\] \[\pm \frac{z_{.005}(\sigma)}{\sqrt n}=E-\mu\] \[\mu\pm \frac{z_{.005}(\sigma)}{\sqrt n}=E\]

OpenStudy (anonymous):

What do I substitute where is what I mean?

OpenStudy (amistre64):

the (+-) part is the spread, or the margin of error, i believe you know n \(\sigma=\sqrt {\hat p(1-\hat p)}\) and the zscore related to 99% about the mean leaves 1%/2 = .005 to look up

OpenStudy (amistre64):

in this case, sigma= sqrt(p^2) = p since p=.5 and 1-p = .5

OpenStudy (anonymous):

I'm not following...

OpenStudy (amistre64):

\[\pm \frac{2.576(.5)}{\sqrt n}=Error_m\]

OpenStudy (anonymous):

I got it know thanks.

OpenStudy (anonymous):

now not Know